OpenStudy (anonymous):

Solve the system for A,B, and C: {A+B+C=0; 7A+4B+C= -26 10A-5B-2C = -47}

OpenStudy (compassionate):

What do yu think you should use, the elimination method or substitution?

OpenStudy (compassionate):

I'll give you a hint, look at + C and -2C

OpenStudy (anonymous):

you can use the matrix method if you know it, its faster and easier, otherwise follow comp. advise

OpenStudy (compassionate):

ISutts, I'm not going to solve it for you, I'm going to work with you and teach you how to solve it. So please reply or I will no longer follow this thread.

OpenStudy (anonymous):

i don;t know either method

OpenStudy (anonymous):

and i have the answers of a = 6, b = -1 and c= -4, but how is what i want to know

OpenStudy (anonymous):

let A+B+C=0.............(i) 7A+4B+C= -26........(ii) 10A-5B-2C = -47..............(iii) we'll use substitution method. from (i) C = 0 -(A+B) = -A-B substitute this in (ii) and (iii) (ii) therefore becomes 7A +4B-A-B = -26 ; it reduces to 6A-3B = -26 (iii) becomes 10A - 5B -2(-A-B) = -47 ; it reduces to 12A - 3B = -47 do you see the common 3B in (ii) and (iii) subtract (ii) from (iii) 12A-6A-3B--3B = -47 -- 26 6A = -21 A = -3.5 B = -15 2/3 C = 19 1/6 Thoat's how to do it and those are these correct answers